# SPFFNet: Strip Perception and Feature Fusion Spatial Pyramid Pooling for Fabric Defect Detection

### v1

@@ IV-A Datasets
Self dataset This dataset was collected and labeled and organized by us. The data mainly comes from the workshop of a fabric factory in Jiangsu Province and public images that can be collected on the Internet. After our collection and organization, the final dataset contains a total of 8,645 fabric defect images, which are classified into five types of defects that are most commonly found in the fabric process: missing stitches, broken holes, stain, broken seam, and broken stitches. The dataset is divided into a training set and a test set in a ratio of 2:8. In addition, this paper also uses some image data enhancement methods, such as rotation, translation, scaling, and flipping, to expand the dataset and generate more samples, thereby improving the generalization ability of the model and reducing the risk of overfitting.
IV-B Evaluation Metrics
The mAP (Mean Average Precision) is a widely used evaluation metric in object detection and information retrieval tasks, providing a comprehensive view of a model’s performance by evaluating precision across different levels of recall. mAP is computed by averaging the Average Precision (AP) for each class, which is the area under the precision-recall curve for that class, and then averaging these values across all classes. The formula for AP is given by:
[…]
where $P(r)$ denotes precision at a given recall level $r$. The final mAP score is calculated as:
$mAP=\frac{1}{N}\sum_{i=1}^{N}AP_{i}.$ (7)
where $N$ is the number of classes and $AP_{i}$ is the Average Precision for class $i$. GFLOPs (Giga Floating Point Operations) is a metric used to measure computational complexity, representing the number of floating-point operations a model performs per second, typically expressed in billions. A lower GFLOPs value indicates better computational efficiency and faster inference times, as fewer operations are required to process the same task. Params (Parameters) refers to the total number of parameters in a model, which reflects its complexity and memory footprint. A lower number of parameters often suggests more memory-efficient models, which can lead to better scalability and less resource consumption. Together, these metrics provide a holistic assessment of a model’s performance, efficiency, and resource utilization, helping to balance the trade-offs between computational power, memory usage, and model accuracy.
mAP@0.5/Method Knot Triple Wire Coarse Pick Broken Spandex Warp Knot Weft Shrink Hole Stain All GFLOPs Params YOLOv5 61.6 77.2 59.4 76.8 45.8 46.9 83.7 43.5 61.9 5.8 2183224 YOLOv6 58.1 78 52.8 68.1 47.1 30.8 82.3 44.5 57.7 11.5 4155816 YOLOv8 65.9 78.8 60.5 76.3 51.3 40.1 81.6 59.9 64.3 6.8 2685928 YOLOv9t 65.4 80.4 59.8 71.8 52.6 46.9 83.3 62.9 65.4 6.4 1731384 YOLOv9s 66 82 54.3 76.6 54.4 46.7 79.7 64.4 65.5 22.1 6196744 YOLOv10n 59.3 77.4 57.7 69.4 41.5 39.2 81.7 57.7 60.5 8.2 2697536 YOLOv11n 64.4 80 64.3 76.1 48.1 43.7 80.5 62.9 65 6.3 2583712 Ours 64.5 80.5 63.5 74.6 49 43.9 83.7 66.4 65.8 6.8 2858951
IV-C Implementation Details
In all experiments, the model size selected for the YOLO series of models is normal. The batch size for training the model is 32, and the input size of the image is 640. Because the dataset has a large number of samples and may contain noisy data, in order to avoid local optima and obtain better model performance, the optimizer selects Stochastic ic gradient descent (SGD), with an initial learning rate of 0.01 and momentum of 0.937. To compare the performance of models of different sizes, the experiment uniformly sets the patience to 20, which is the number of epochs that the training is allowed to continue without improving the accuracy of the model on the validation set.
IV-D Comparison with State-of-the-art Methods
We compared the proposed method with six state-of-the-art methods, including YOLOv5 [20], YOLOv6 [33], YOLOv8 [22], YOLOv9t [34], YOLOv9s [34], and YOLOv10n [35].
IV-D1 Comparisons on Tianchi fabric dataset
@@
IV-D2 Comparisons on Self dataset
@@
TABLE II: Comparison of the performance of the proposed improved model with multiple SOTA on the Self dataset
mAP@0.5/Method Missing Stitches Broken Holes Stain Broken Seam Broken Stitches All GFLOPs Params YOLOv5 85.4 73.4 99.5 80.2 75.9 82.9 5.8 2182639 YOLOv6 83 68.9 99.5 80.2 55.5 77.4 11.5 4155519 YOLOv8 93.9 78.2 99.5 82 88.1 88.3 6.8 2685343 YOLOv9t 89.1 76.3 99.5 82.1 85.8 86.5 6.4 1730799 YOLOv9s 91.7 80.2 99.5 81.2 91.8 88.9 22.1 6195583 YOLOv10n 89.5 76.8 99.5 78.6 85.9 86.1 8.3 2696336 YOLOv11n 93.1 79.4 99.5 83.8 89.3 89 6.3 2583127 Ours 95.3 83.5 99.5 81.1 93.5 90.6 6.8 2858951
Figure 4: Comparison visualized by heat maps
IV-E Ablation Studies and Analysis
The comparison results in Tables I and II show that the proposed improved model is superior to many state-of-the-art single-stage detection methods. Next, a comprehensive analysis of the proposed improved model will be performed by testing it on the dataset we created to explore the logical basis for its superiority. As shown in Table III, the model containing the SPM, SE-SPPF, and FECIoU modules has the highest detection accuracy, with an mAP of 90.6%. This is an improvement over the baseline model, which has an mAP of 89%. The baseline model does not include these modules, and its computational cost is 6.3 GFLOPs and the number of parameters is 2.58 million. After the SPM module is introduced into the model, the detection accuracy is improved to 89.6%, the computational cost is slightly increased to 6.6 GFLOPs, and the number of parameters is slightly increased to 2.61 million. This indicates that the SPM module has the effect of enhancing the extraction of features for strip defects in the dataset. Similarly, when the SE-SPPF module is added alone, the detection accuracy is 89.6%, and GFLOPs (6.6) and parameters (2.89 million) increase slightly, which indicates that SE-SPPF also plays a key role in defect feature extraction by better fusing channel and spatial features. When both the SPM and SE-SPPF modules were included, the mAP was further improved to 90.3%, with a computational cost of 6.8 GFLOPs and a parameter count of 2.86 million. This indicates that the combination of these modules enhances feature extraction capabilities without a significant increase in computational cost. Finally, when the three components SPM, SE-SPPF and FECIoU are integrated, the model achieves the highest mAP of 90.6%, with a slight increase in computational cost (6.8 GFLOPs) and 2.86 million parameters. This shows the synergistic effect of these modules, as they work together to improve the accuracy of the model while maintaining a reasonable balance of computational efficiency..
TABLE III: Results of ablation experiments on self datasets
SPM SE-SPPF FECIoU mAP@0.5/GFLOPs Params
@@ IV-E Ablation Studies and Analysis
IV-F Visualization
@@ V Conclusion
This paper introduced an enhanced fabric defect detection model built upon YOLOv11. To improve the model’s ability to capture and extract features of stripe defects, a SPM was designed and incorporated. Additionally, the SPPF was enhanced, and a novel Squeeze-and-Excitation Spatial Pyramid Pooling Fast (SE-SPPF) was proposed to strengthen the model’s capacity to differentiate backgrounds and extract defect features. Moreover, FECIoU was proposed, an adaptive-weight version of the CIoU, to mitigate the effects of significant scale differences between target boxes. SPM utilized multi-scale convolution to effectively capture features at various scales within the feature map, while its dense connection structure enhanced the accuracy and efficiency of feature sharing, leading to an overall improvement in the model’s accuracy and speed. SE-SPPF combined weighted channel feature maps with spatial pyramid pooling, ensuring the comprehensive integration of both spatial and channel information, which further boosted the model’s ability to extract complex features. FECIoU applied focal loss to adjust the weights of hard-to-detect instances during training, addressing class imbalance issues and ultimately improving the overall detection performance. In conclusion, the proposed model outperformed other state-of-the-art methods, achieving an increase in mAP of 0.8-8.1% on the Tianchi dataset and 1.6-13.2% on our custom dataset. However, there are still some limitations to the current work. For example, the types of fabric defects currently studied are too few, and the types of defects in actual production are far more than those in the current dataset. The performance of the model on new defect types needs to be further explored. There is also a defect of color error in the fabric industry, which changes with the color of the fabric and poses a new challenge to the model.

### v3

@@ IV-A2 Produced dataset
This dataset was collected and labeled and organized by us. The data mainly comes from the workshop of a fabric factory in Jiangsu Province and public images that can be collected on the Internet. After our collection and organization, the final dataset contains a total of 8,645 fabric defect images, which are classified into five types of defects that are most commonly found in the fabric process: missing stitches, broken holes, stain, broken seam, and broken stitches. In addition, this paper also uses some image data enhancement methods, such as rotation, translation, scaling, and flipping, to expand the dataset and generate more samples, thereby improving the generalization ability of the model and reducing the risk of over-fitting.
IV-B Implementation Details
All experiments were conducted on an NVIDIA RTX 4090D GPU, with the YOLO series models configured to use their standard (“normal”) size variants. The models were trained with a batch size of 32 and an input resolution of 640 × 640. Given the large scale of the dataset and the potential presence of noisy samples, Stochastic Gradient Descent (SGD) was adopted as the optimizer to enhance convergence stability and mitigate the risk of local minima, with an initial learning rate of 0.01 and momentum of 0.937. To ensure fair comparison across models of different sizes, the early stopping patience was uniformly set to 20 epochs, allowing training to continue for up to 20 epochs without improvement in validation accuracy before termination.
IV-C Comparison with State-of-the-art Methods
We compare the proposed SPFFNet with six state-of-the-art object detection models, including YOLOv5 [jocher_2020_ultralyticsyolov5], YOLOv6 [li_2022_yolov6], YOLOv8 [jocher_2023_yolov8], YOLOv9-t [wang_2024_yolov9], YOLOv9-s [wang_2024_yolov9], and YOLOv10-n [wang_2024_yolov10], to comprehensively assess detection accuracy and efficiency under consistent experimental conditions.
IV-C1 Comparisons on Tianchi fabric dataset
@@
IV-C2 Comparisons on produced dataset
@@
TABLE III: Results of ablation experiments on produced datasets
SPM SE-SPPF FECIoU mAP@0.5/% GFLOPs Params
@@
IV-D Ablation Studies and Analysis
Tables I and II demonstrate that the proposed model consistently outperforms several state-of-the-art single-stage detectors. To further substantiate its effectiveness, an ablation study was conducted on the custom dataset (Table III). Integrating the SPM, SE-SPPF, and FECIoU modules yields the best performance, achieving 90.6% mAP with only a marginal increase in computation (6.8 vs. 6.3 GFLOPs). Specifically, SPM enhances strip-oriented feature perception, while SE-SPPF strengthens spatial–channel interactions; both contribute notable accuracy improvements with negligible overhead. Their synergistic combination demonstrates strong complementarity, and the inclusion of FECIoU further refines localization, resulting in the highest overall accuracy. These findings confirm that the proposed components effectively boost detection capability while maintaining computational efficiency, underscoring the robustness and practicality of SPFFNet for real-world fabric defect detection.
Figure 4: Comparison visualized by heat maps
IV-E Visualization
@@ V Conclusion
In this paper, we propose SPFFNet, an enhanced fabric defect detection framework built upon YOLOv11, which integrates the Strip Perception Module (SPM), Squeeze-and-Excitation Spatial Pyramid Pooling Fast (SE-SPPF), and Focal Enhanced Complete IoU (FECIoU) loss to improve feature representation, background discrimination, and localization precision. Extensive experiments on the Tianchi and custom datasets demonstrate that SPFFNet achieves consistent gains over state-of-the-art approaches, confirming its effectiveness for complex industrial inspection scenarios.
However, the current model is still limited by its reliance on RGB imagery and a relatively narrow range of defect categories, which may restrict its generalization to diverse textile materials and illumination conditions. Future work will focus on enhancing the model’s robustness to color variations and unseen defect patterns through cross-domain learning and spectral feature integration, as well as improving its efficiency and adaptability for real-time deployment in large-scale manufacturing environments.